OpenStudy (anonymous):

int from 1 to infinity of (1/ sqrt(x+4))

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty} 1/(\sqrt[?]{x+4})\]

OpenStudy (anonymous):

try \(u=x+4\)

OpenStudy (anonymous):

\[\int\limits_{1}^{t} 1/(\sqrt{x+4}) dx\]

OpenStudy (anonymous):

oops forgot lim as t approaches infinity

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}du/\sqrt{u}\]

OpenStudy (anonymous):

like @satellite73 said

OpenStudy (anonymous):

you get \[2\sqrt{x+4}|_1^t\] now fairly clearly this will not converge as \(t\to\infty\)

OpenStudy (anonymous):

with a little experience you will be able to say "no way" before even starting. if you want to integrate to infinity, the power in the denominator has to be bigger than one, and you only have a power of \(\frac{1}{2}\)

OpenStudy (anonymous):

\[\lim t \rightarrow \infty \int\limits_{1}^{t} (1/\sqrt{u}) du\]

OpenStudy (anonymous):

sorry i'm reading the comments now

OpenStudy (anonymous):

\[\lim_{t\to \infty}\sqrt{t+4}\] is not finite

OpenStudy (anonymous):

why doesn't it converge when 2 sqrt(1+4) = 2 sqrt(5)

OpenStudy (anonymous):

oh you can plug in1 alright, that is no problem, one is a number

OpenStudy (anonymous):

but you also have to plug in \(t\) and then take the limit as \(t\to\infty\)

OpenStudy (anonymous):

that is infinite, so no integral

OpenStudy (anonymous):

which is an infinitly big number...true

OpenStudy (anonymous):

so I should be able to look at the equation and see that the denominator becomes infinitely big and assume it doesn't converge?

OpenStudy (anonymous):

one more question plz

OpenStudy (anonymous):

so the equation \[\sum_{1}^{\infty} 1/n^5\] convergest because the x in the numerator?

OpenStudy (anonymous):

converges i mean

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty} 1/x^5 dx\]

OpenStudy (anonymous):

converges because \(5>1\)